# Monte Carlo Tree Diffusion with Multiple Experts for Protein Design

### v1

@@ Abstract
The goal of protein design is to generate amino acid sequences that fold into functional structures with desired properties. Prior methods combining autoregressive language models with Monte Carlo Tree Search (MCTS) struggle with long-range dependencies and suffer from an impractically large search space. We propose MCTD-ME, Monte Carlo Tree Diffusion with Multiple Experts, which integrates masked diffusion models with tree search to enable multi-token planning and efficient exploration. Unlike autoregressive planners, MCTD-ME uses biophysical-fidelity-enhanced diffusion denoising as the rollout engine, jointly revising multiple positions and scaling to large sequence spaces. It further leverages experts of varying capacities to enrich exploration, guided by a pLDDT-based masking schedule that targets low-confidence regions while preserving reliable residues. We propose a novel multi-expert selection rule ($\text{P}\mathcal{H}\text{-UCT}\text{-ME}$) extends predictive-entropy UCT to expert ensembles. On the inverse folding task (CAMEO and PDB benchmarks), MCTD-ME outperforms single-expert and unguided baselines in both sequence recovery (AAR) and structural similarity (scTM), with gains increasing for longer proteins and benefiting from multi-expert guidance. More generally, the framework is model-agnostic and applicable beyond inverse folding, including de novo protein engineering and multi-objective molecular generation.
@@ 1 Introduction
Large language models (LLMs such as GPT [18, 19, 22, 23, 24]) have recently been paired with MCTS to improve search in complex generation tasks. For instance, GPT-guided MCTS has enhanced search efficiency in symbolic regression [15], while value-guided MCTS decoding (PPO-MCTS [16]) improved the preferability of generated text over standard RLHF outputs. However, GPT-based MCTS approaches face notable pitfalls. Due to the autoregressive, token-by-token generation, even a single token error can derail an entire solution, and models often get stuck in suboptimal trajectories because their prior training biases lead to unbalanced exploration-exploitation. In other words, maintaining long-range coherence is challenging and it’s difficult to revise parts of a candidate without regenerating the whole sequence.
Beyond these GPT-specific pitfalls, integrating MCTS into large-scale generative frameworks—particularly for biological design—faces major challenges as follows: First, unlike classical planning environments, we lack a tractable forward model to evaluate partially generated sequences. While recent approaches like Diffuser [9] show how diffusion models can operate as planners without explicit forward simulators, standard diffusion decoding is still largely open-loop. This limits its ability to adaptively revise candidates during generation. Second, real-world biological design tasks are intrinsically multi-objective. Protein design, for instance, requires satisfying multiple criteria, from structural compatibility to folding efficiency and evolutionary plausibility. Third, GPT-based generators make it inefficient to apply MCTS, especially for long-sequence tasks, resulting in unnecessarily large tree searches in both breadth and depth and inevitably exploring unpromising states. Finally, relying on a single GPT for exploration limits biophysical-fidelity, diversity, and constrains the search space due to biases from the pretraining dataset.
@@
In this paper, we propose Monte Carlo Tree Diffusion with Multiple Experts (MCTD-ME), a novel planning framework that integrates diffusion-based generation with an MCTS search guided by a diverse ensemble of domain experts models. Motivated by the challenges of protein sequence design—where search spaces are vastly larger than small-molecule generation—we leverage diffusion models for their ability to efficiently denoise multiple positions in parallel, and use expert models to narrow the search space by injecting biologically grounded priors. Each node in our tree represents a partially denoised sequence (a partial plan), and expansion corresponds to one step of noise reduction. This structure imbues the diffusion process with MCTS’s adaptive lookahead: evaluating, pruning, and refining paths to focus on promising candidates. In addition, we adopt an imitation-style rollout approach, where each expert proposes candidate completions based on the current partial sequence, thereby increasing diversity in the exploration space. These trajectories guide selection and backpropagation through the tree. Disagreement among experts encourages exploration; consensus promotes exploitation—enabling informed multi-objective planning at test time. We defer additional details on related work to Appendix 2.
@@
Summary of contributions: To our knowledge, this is the first work to incorporate an ensemble of expert into a diffusion-based planning framework, enabling multi-experts Monte Carlo tree search for complex generative problems. By unifying diffusion models with a multi-critic evaluation module and a multi-expert proposal module within MCTS, we introduce a general test-time strategy to steer generation toward goals that single decoders or static samplers struggle with—especially in large, structured domains like protein sequence design, where generation spaces are vast and require both parallel decoding and multi-perspective exploration. Central to our design are (i) a pLDDT-guided masking policy that focuses edits on structurally uncertain regions to improve biophysical fidelity, and (ii) a multi-expert selection mechanism that arbitrates among experts using cached uncertainty statistics, yielding stable and efficient search.
@@
    A tree-structured diffusion process that expands and evaluates partial sequences through iterative biophysical-fidelity-enhanced denoising, achieved by masking low-pLDDT positions and guided by a multi-expert rollout mechanism, where each expert proposes candidate completions conditioned on the partial sequence, thereby enabling more diverse and meaningful exploration across a vast search space.
@@
    A novel $\text{P}\mathcal{H}\text{-UCT}\text{-ME}$ selection algorithm, an entropy-aware selection rule, extending the classic UCT formula with exploration bonuses based on expert disagreement (e.g., score variance or entropy), inspired by entropy-guided UCB from ERP [21].
Third, we demonstrate empirical gains on long-horizon generation tasks—specifically inverse protein folding for just one case example. Using the CAMEO 2022 benchmark [10] and a PDB date-split benchmark [4], we show that our framework consistently improves sequence recovery (AAR) and structural similarity (scTM) over strong DPLM-2 baselines [31]. Our approach yields higher-quality, more diverse solutions, and these improvements grow with larger search budgets—highlighting the scalability and deliberative nature of our framework. We observe that both pLDDT-guided masking and multi-expert selection contribute complementary gains, with larger benefits on longer proteins where uncertainty is spatially concentrated. Overall, this work advances the state-of-the-art in test-time guided generative planning, opening up promising directions for applications in drug discovery, program synthesis, and beyond.
@@ 2 Related works
We review related work in protein diffusion models, MCTS in generative modeling and molecule design, discrete diffusion for planning and generation, and multi-objective and expert-guided Generation.
Protein and Drug Diffusion Models
Recent advances apply diffusion models to protein design across both structure and sequence domains. RFdiffusion [32] introduced a guided 3D diffusion model for protein backbones, achieving state-of-the-art results. On the sequence side, DPLM [30] and its extension DPLM-2 [31] demonstrated that discrete diffusion pretraining on large-scale protein data enables high-quality sequence generation and versatile conditioning (e.g., inverse folding, property steering). Most recently, DiffGui [8] proposed a target-aware, E(3)-equivariant diffusion framework for joint atom and bond generation, incorporating multi-property guidance (e.g., binding affinity, drug-likeness) to generate realistic, high-affinity 3D molecules. However, these diffusion approaches generally rely on one-shot or open-loop decoding, making it difficult to incorporate adaptive search or revision during generation—highlighting the need for planning methods such as MCTS.
Monte Carlo Tree Search in Generative Modeling
Monte Carlo Tree Search (MCTS) has improved generative decoding in tasks requiring complex reasoning or program synthesis. In code generation, AlphaCode [14] used a brute-force generate-and-test method by sampling large program sets and selecting those that passed tests. More structured methods followed: RethinkMCTS [13] performs MCTS over LLM reasoning steps, using execution feedback to refine intermediate thoughts. Tree-of-Thoughts [33] frames reasoning as a tree of self-evaluated partial solutions expanded via classical search (DFS/BFS), boosting performance on puzzles and planning. PG-TD [34] simulates lookahead in code generation by executing candidate programs during decoding. Across domains, MCTS-style planning consistently enhances generation quality and reliability. Yet, most of these approaches are tied to autoregressive GPT-style models, which remain token-by-token and struggle with long-range dependencies—suggesting that discrete diffusion could provide a more flexible backbone for planning.
Discrete Diffusion for Planning and Generation
While originally developed for continuous domains, diffusion models have been successfully adapted to discrete sequence generation and planning. D3PM [1] introduced absorbing-state (masking) noise for discrete diffusion, achieving competitive text generation. Follow-ups like [27] further closed the performance gap with autoregressive models. In planning, Diffuser [9] modeled offline RL as trajectory denoising, generating full state-action paths guided by value functions. To enable adaptive inference, Monte Carlo Tree Diffusion (MCTD) [2] reimagines diffusion as tree search, branching multiple denoising outcomes per step and scoring partial trajectories via heuristics. In protein design, ProtInvTree [17] similarly integrates reward-guided search with jumpy denoising for efficient inverse folding. Still, existing methods optimize toward a single diffusion generator, which limits their ability to effective search space exploration in biological and chemical design.
Multi-Experts Guided Generation
Multi-expert systems leverage several generative models or “experts” in tandem, allowing each to contribute from its area of strength. This idea has been explored in reinforcement learning and bandit settings, where algorithms learn from multiple expert policies or oracles to improve decision-making [20]. In generative modeling, mixture-of-experts strategies similarly combine specialized models to better cover complex data distributions. For example, recent diffusion-based image generators use multiple specialized denoising models across different noise levels [5] or even heterogeneous model architectures [12] to boost output quality. However, these multi-expert approaches remain largely static and one-shot during inference, without an adaptive search process. To our knowledge, no prior diffusion model integrates multiple experts into a planning loop for generative design. By incorporating a team of expert generators within an MCTS-driven diffusion framework, our work enables dynamic collaboration among experts to satisfy multiple design criteria and explore a broader solution space—an innovative combination that extends diffusion-based generation to more complex, multi-objective scenarios.
@@ 3 Preliminaries
We describe Diffusion-based generative model, MCTS, and protein design along with their mathematical notation in the subsequent sections and unify them under the Markov decision processes framework.
@@
We formulate generative planning as a finite-horizon MDP $\mathcal{M}=\langle\mathcal{S},\mathcal{A},H,P,R\rangle$ [26], where $\mathcal{S}$ is the set of states (e.g., partial or completed sequences), $\mathcal{A}$ is the action space (e.g., denoising steps or token insertions), and $P:\mathcal{S}\times\mathcal{A}\to\mathcal{S}$ defines stochastic transitions. The reward function $R:\mathcal{S}\times\mathcal{A}\to\mathbb{R}$ provides a scalar evaluation, often defined only on terminal states. The policy $\pi:\mathcal{S}\to\Delta(\mathcal{A})$ defines the action distribution at each state, such as a denoising model in diffusion or a next-token model in autoregressive generation. An episode corresponds to constructing a sequence $y=(y_{1},...,y_{H})$, with a termination action marking the end of generation. We denote the action-value function $Q^{\pi}(s,a)$ as the expected cumulative reward of taking action $a$ in state $s$ under policy $\pi$, and the value function $V^{\pi}(s)$ as the expected return from $s$.
@@
We briefly review the diffusion framework for sequence generation. Let $y_{0:H}$ denote a token sequence of length $H$. A forward noising process $q(y_{t}\mid y_{t-1})$, for $t=1,\ldots,T$, gradually corrupts the sequence, such that $y_{T}$ becomes maximally noisy (e.g., fully masked for discrete data or Gaussian noise for continuous). A reverse denoising model $p_{\phi}(y_{t-1}\mid y_{t})$ is trained to recover $y_{t-1}$ from $y_{t}$, enabling sampling via iterative denoising from $y_{T}\sim p(y_{T})$ down to $y_{0}\sim p_{\phi}(y_{0})$. Crucially, diffusion models naturally support guided generation. When an auxiliary score function $f(y)$ is available—e.g., a property predictor or constraint critic—generation can be biased toward desirable outputs. One classic approach is classifier guidance, where each reverse step is adjusted in the direction of $\nabla_{y_{t-1}}f(y_{t-1})$. Alternatively, reverse distribution can be reweighted as:
@@
with $\beta$ controlling guidance strength. This formulation generalizes to multiple critics: given $K$ evaluators $C_{0},\dots,C_{K-1}$, each scoring complete sequences, guidance becomes even more natural. Instead of aggregating into a scalar reward, we perform imitation-style rollouts: partial candidates are completed, scored individually by each expert, and used to inform node evaluations during tree search. This preserves the diversity of expert preferences and supports structured planning, which we realize through our Monte Carlo Tree Diffusion (MCTD) framework.
@@
Monte Carlo Tree Search (MCTS) is a planning algorithm that combines tree-based exploration with stochastic simulation to balance exploration and exploitation in large decision spaces. It proceeds in four stages: selection, expansion, simulation, and backpropagation. Starting from the root (current state), the selection phase recursively chooses child nodes using an Upper Confidence Bound (UCT) criterion [11]:
@@
where $Q(s,a)$ is the estimated reward, $N(s)$ and $N(s,a)$ are visit counts, and $c_{p}$ controls exploration. When a leaf node is reached, expansion adds child nodes corresponding to unexplored actions. Then, simulation evaluates these new nodes by executing a rollout policy (e.g., sampling until termination or a predefined depth), producing a scalar reward. Finally, backpropagation propagates the reward up the visited path, updating $N(\cdot)$ and $Q(\cdot)$ statistics. Over repeated iterations, MCTS refines its value estimates and focuses the search on high-reward regions of the state space.
@@ 4 The MCTD-ME Algorithm
Figure 1: Overview of MCTD in diffusion-based protein sequence generation. Nodes are partially denoised sequences expanded via masked diffusion and guided by multi-expert rollouts.
We propose Monte Carlo Tree Diffusion with Multi-Experts (MCTD-ME), which enhances protein generation by combining masked diffusion with an MCTS planner. MCTD-ME structures diffusion as a tree, refining partially masked sequences to preserve scaffold integrity while systematically exploring uncertain regions. Search quality and diversity are further improved by incorporating multiple diffusion experts for diverse refinements and by using pLDDT-guided masking to target low-confidence residues. These design choices couple the planning strength of tree search with expert-aware diversification, yielding a principled framework for balancing exploration and exploitation in protein modeling. We present the algorithm in Algorithm 1, with each rollout consisting of four key steps.
@@
To navigate the space of protein sequences under the reverse masked diffusion framework, we define a tree where each node corresponds to a complete sequence (with no masked tokens) at a particular reverse step, and edges represent one reverse transition obtained by applying masked discrete diffusion to a low-confidence subset of tokens. At each iteration, we employ a novel selection algorithm, $\text{P}\mathcal{H}\text{-UCT}\text{-ME}$—an entropy-augmented variant of UCT—designed to identify the most promising path, where ME denotes Multiple Experts. Specifically, from a parent node $s_{t}$ we choose the transition (i.e., reverse move to $s_{t-1}$ via a particular mask set) that maximizes the following objective:
$\text{P}\mathcal{H}\text{-UCT}\text{-ME}(s_{t})\;=\;\mathop{\rm arg\,max}_{a\in\mathcal{A}}\;\text{P}\mathcal{H}\text{-UCB}\text{-ME}\!\left((s_{t},a)\right),$ (2)
@@
$\text{P}\mathcal{H}\text{-UCB}\text{-ME}\!\left((s_{t},a)\right)$ $=Q(s_{t},a)+c_{p}\,\frac{\sqrt{\log N(s_{t})}}{1+N(s_{t},a)}\cdot\left({\underbrace{w_{\mathrm{ent}}\,\mathcal{U}_{\text{ent}}(s_{t},a)\;+\;w_{\mathrm{div}}\,\mathcal{U}_{\text{div}}(s_{t},a)}_{\text{$\text{P}\mathcal{H}$\text{-ME} bonus (cached at expansion)}}}\right).$ (3)
Here, $Q(s_{t},a)$ is the current value estimate, $N(\cdot)$ are visit counts, $c_{p}>0$ controls exploration, and $w_{\mathrm{ent}},w_{\mathrm{div}}\!\geq\!0$ weight the predictive-uncertainty and diversity bonuses, respectively. In our diffusion setting, the action prior is given by the reverse diffusion model $p_{\phi}(s_{t-1}\!\mid\!s_{t})$ over masked edits, but—consistent with our implementation—we do not recompute its entropy during selection; instead we cache uncertainty statistics for each expanded child and reuse them during subsequent selections.
Multi-expert PH bonus. Let $\mathcal{M}$ be the mask set chosen at $s_{t}$ (e.g., via pLDDT), and let $\{p_{\phi}^{(e)}\}_{e=1}^{E}$ denote the $E$ expert conditional distributions (softmax of logits) used during expansion. For the candidate child $s_{t-1}$ obtained by action $a$, we define the ensemble-based uncertainty as
$\mathcal{U}_{\text{ent}}(s_{t},a)=\mathcal{H}\!\left(\frac{1}{E}\sum_{e=1}^{E}p_{\phi}^{(e)}(\,\cdot\,\mid s_{t-1},\mathcal{M})\right)-\frac{1}{E}\sum_{e=1}^{E}\,\mathcal{H}\!\left(p_{\phi}^{(e)}(\,\cdot\,\mid s_{t-1},\mathcal{M})\right).$ (4)
Here $\mathcal{H}[p]=-\sum_{a}p(a)\log p(a)$ denotes Shannon entropy. The first term is the entropy of the ensemble-averaged distribution $\bar{p}=\tfrac{1}{E}\sum_{e}p^{(e)}$, while the second is the average entropy of the individual experts. Their difference equals the mutual information (BALD score [7]), or Jensen–Shannon–style divergence between experts. Intuitively, it isolates epistemic uncertainty: if experts disagree, $\mathcal{H}[\bar{p}]$ is large while the average $\mathcal{H}[p^{(e)}]$ remains small, giving a positive signal. If all experts are merely noisy but agree, the two terms cancel.
@@
$\mathcal{U}_{\text{div}}(s_{t},a)$ $=\frac{1}{L}\sum_{i=1}^{L}\,\mathbf{1}\{y^{\text{child}}_{i}\neq y^{\text{parent}}_{i}\},$ (5)
For our inverse folding case studies, this simply counts the fraction of amino acids that differ. Both $\mathcal{U}_{\text{ent}}$ and $\mathcal{U}_{\text{div}}$ are computed once at expansion and stored for reuse during selection.

### v2

@@ Abstract
The goal of protein design is to generate amino acid sequences that fold into functional structures with desired properties. Prior methods combining autoregressive language models with Monte Carlo Tree Search (MCTS) struggle with long-range dependencies and suffer from an impractically large search space. We propose MCTD-ME, Monte Carlo Tree Diffusion with Multiple Experts, which integrates masked diffusion models with tree search to enable multi-token planning and efficient exploration under the guidance of multiple experts. Unlike autoregressive planners, MCTD-ME uses biophysical-fidelity-enhanced diffusion denoising as the rollout engine, jointly revising multiple positions and scaling to large sequence spaces. It further leverages experts of varying capacities to enrich exploration, guided by a pLDDT-based masking schedule that targets low-confidence regions while preserving reliable residues. We propose a novel multi-expert selection rule ($\text{P}\mathcal{H}\text{-UCT}\text{-ME}$) extends Shannon-entropy-based UCT to expert ensembles with mutual information. MCTD-ME achieves superior performance on the CAMEO and PDB benchmarks, excelling in protein design tasks such as inverse folding, folding, and conditional design challenges like motif scaffolding on lead optimization tasks. Our framework is model-agnostic, plug-and-play, and extensible to de novo protein engineering and beyond.
Machine Learning, ICML
[…]
$\Statex\ALG@thistlm\triangleright #1 \algnewcommand\IndentLineComment$$\Statex\ALG@tlm\triangleright #1$
@@ 1 Introduction
Large language models (LLMs such as GPT [25, 26, 27]) have recently been paired with MCTS to improve search in complex generation tasks. For instance, GPT-guided MCTS has enhanced search efficiency in symbolic regression [17], while value-guided MCTS decoding (PPO-MCTS [19]) improved the preferability of generated text over standard RLHF outputs. However, GPT-based MCTS approaches face notable pitfalls. Due to autoregressive, token-by-token generation, even a single token error can derail an entire solution, and models often get stuck in suboptimal trajectories because their prior training biases lead to unbalanced exploration-exploitation. In other words, maintaining long-range coherence is challenging and it’s difficult to revise parts of a candidate without regenerating the whole sequence.
Beyond these GPT-specific pitfalls, applying MCTS to biological sequence design raises core problems: (i) no reliable partial-state evaluator (i.e., no tractable forward model) to score or prune partial sequences; (ii) multi-objective constraints (e.g., structural compatibility, folding efficiency, evolutionary plausibility); (iii) poor scaling to long sequences under autoregressive generation, making tree search expensive and prone to exploring unpromising states; and (iv) limited diversity/biophysical fidelity when exploration relies on a single pretrained model. While Diffuser [11] shows diffusion models can act as planners without explicit forward simulators, diffusion decoding is still largely open-loop, offering limited opportunities for test-time correction.
@@
In this paper, we propose Monte Carlo Tree Diffusion with Multiple Experts (MCTD-ME), a novel planning framework that integrates diffusion-based generation with an MCTS search guided by a diverse ensemble of domain experts models. Motivated by the challenges of protein sequence design—where search spaces are vastly larger than small-molecule generation—we leverage diffusion models for their ability to efficiently denoise multiple positions in parallel, and use expert models to narrow the search space by injecting biologically grounded priors. Each node in our tree represents a partially denoised sequence (a partial plan), and expansion corresponds to one step of noise reduction. This structure imbues the diffusion process with MCTS’s adaptive lookahead: evaluating, pruning, and refining paths to focus on promising candidates. In addition, we adopt an imitation-style rollout approach, where each expert proposes candidate completions based on the current partial sequence, thereby increasing diversity in the exploration space. These trajectories guide selection and backpropagation through the tree. Disagreement among experts encourages exploration; consensus promotes exploitation—enabling informed multi-objective planning at test time.
@@
Summary of contributions: To our knowledge, this is the first work to incorporate an ensemble of experts into a diffusion-based planning framework, enabling multi-expert Monte Carlo tree search for complex generative problems. By unifying diffusion models with a multi-critic evaluation module and a multi-expert proposal module within MCTS, we introduce a general inference-time strategy to steer generation toward goals that single decoders or static samplers struggle with—especially in large, structured domains like protein sequence design, where generation spaces are vast and require both parallel decoding and multi-perspective biophysical-fidelity-aware exploration.
@@
    A tree-structured diffusion process for partial sequences, where each node represents a masked-infill state and expansion corresponds to a denoising rollout. We improve biophysical fidelity via a pLDDT-guided masking policy that focuses edits on structurally uncertain regions, and we enable diverse, goal-directed exploration using a multi-expert rollout mechanism, where each expert proposes candidate completions conditioned on the partial sequence.
@@
    A novel $\text{P}\mathcal{H}\text{-UCT}\text{-ME}$ selection rule for multi-expert MCTS, which arbitrates among experts using cached uncertainty statistics for stable and efficient search, extending the classic UCT formula with exploration bonuses based on expert disagreement (e.g., score variance or entropy), inspired by entropy-guided UCB [23].
Third, we demonstrate empirical gains on long-horizon generation tasks—folding, inverse folding, and motif scaffolding—on CAMEO 2022 [12] and a PDB date-split benchmark [5]. In the lead-optimization regime (equal test-time compute), our planner consistently improves over strong baselines such as ESMFold [18], DPLM-2 [37], and ProteinMPNN [7] (we report absolute baseline/final values and $\Delta$). We obtain lower RMSD and higher TMscore in folding, higher AAR and scTM in inverse folding, and improved motif preservation in scaffolding; moreover, we outperform single-expert ablations of RFDiffusion [39], FoldFlow [4], and ProteinA [9], with gains amplifying on longer proteins and more challenging scaffolds where uncertainty is concentrated. Beyond lead optimization, we additionally evaluate a de novo design setting (all-mask root) in Appendix A.8.
Our results highlight three practical advantages: (i) pLDDT-guided masking concentrates compute on low-confidence regions, improving efficiency and stability; (ii) multi-expert selection leverages complementary pretrained models, yielding robust gains even when individual experts underperform; and (iii) quality improves smoothly with larger search budgets, enabling controllable compute–performance trade-offs. Importantly, the method is model-agnostic and modular: it can integrate any set of pretrained experts and black-box critics at test time without retraining, making it broadly applicable across domains. Overall, this work advances the state-of-the-art in test-time guided generative planning, opening up promising directions for applications in drug discovery, program synthesis, and beyond.
@@ 2 Related works
Recent work shows diffusion models are promising for biomolecular design, including structure-based backbone generation (RFdiffusion [38]), discrete sequence generation with versatile conditioning (DPLM and DPLM-2 [36, 37]), and property-guided atom–bond generation (DiffGui [10]); however, these methods are typically one-shot and lack test-time refinement. In parallel, tree search has improved generation for code and reasoning (AlphaCode [16], RethinkMCTS [15], Tree-of-Thoughts [40], PG-TD [43]) and more recently System-2-style planning for LMs [2], but largely targets autoregressive GPT-style models rather than diffusion.
Discrete diffusion supports generation via absorbing-state noise (D3PM [1]) and masked-diffusion LMs simplifying denoising [32, 31], plus planning (Diffuser [11]), inverse folding (ProtInvTree [20]), and peptide optimization (PepTune [34]). Closest, masked-diffusion has also been used for planning in diffusion spaces (e.g., path planning) [31, 32], but these methods plan over trajectories, whereas our setting is protein sequence design with biophysical critics and pLDDT-guided edit control. While mixture-of-experts strategies have been explored in RL and diffusion-based image generation [8, 14, 22], they are static at inference. Our approach integrates multiple specialized expert generators within an MCTS-driven diffusion framework for dynamic collaboration, adaptive search, and multi-objective biomolecular design. (Additional details are provided in Appendix A.6.)
@@ 3 Preliminaries
We formulate generative planning as a finite-horizon MDP $\mathcal{MDP}=\langle\mathcal{S},\mathcal{A},H,P,R\rangle$ [30], where $\mathcal{S}$ is the set of states (e.g., partial or completed sequences), $\mathcal{A}$ is the action space (e.g., denoising steps or token insertions), and $P:\mathcal{S}\times\mathcal{A}\to\mathcal{S}$ defines stochastic transitions. The reward $R:\mathcal{S}\times\mathcal{A}\to\mathbb{R}$ is often defined only on terminal states, and the policy $\pi:\mathcal{S}\to\Delta(\mathcal{A})$ specifies an action distribution (e.g., a diffusion denoiser or autoregressive next-token model). An episode constructs a length-$H$ sequence $y=(y_{1},...,y_{H})$. We denote the action-value function $Q^{\pi}(s,a)$ and value function $V^{\pi}(s)$.
@@
We review diffusion for sequence generation. Let $y_{0:H}$ denote an $H$-token sequence. Forward noising $q(y_{t}\mid y_{t-1})$ for $t=1,\ldots,T$ corrupts the sequence until $y_{T}$ is maximally noisy (e.g., fully masked). A reverse model $p_{\phi}(y_{t-1}\mid y_{t})$ is trained to denoise, enabling sampling by iterating from $y_{T}\sim p(y_{T})$ to $y_{0}\sim p_{\phi}(y_{0})$.Crucially, diffusion models naturally support guided generation. When an auxiliary score function $f(y)$ is available—e.g., a property predictor or constraint critic—generation can be biased toward desirable outputs. One classic approach is classifier guidance, where each reverse step is adjusted in the direction of $\nabla_{y_{t-1}}f(y_{t-1})$. Alternatively, reverse distribution can be reweighted as:
@@
with $\beta$ controlling guidance strength. We extend this to multiple experts: $K$ evaluators $\{C_{k}\}_{k=1}^{K}$ score complete sequences proposed by multiple experts $\{\pi^{e}\}_{e=1}^{E}$ .
@@
Monte Carlo Tree Search (MCTS) balances exploration and exploitation via tree search and simulation. It iterates selection, expansion, evaluation, and backpropagation. Selection uses UCT [13]:
@@
where $Q(s,a)$ is the estimated reward, $N(s)$ and $N(s,a)$ are visit counts, and $c_{p}$ controls exploration. Expansion adds children, evaluation runs a rollout to obtain a reward, and backpropagation updates $N(\cdot)$ and $Q(\cdot)$. Repetition concentrates search on high-reward regions.
@@ 4 The MCTD-ME Algorithm
Figure 1: Overview of MCTD-ME in diffusion-based protein sequence generation. Nodes are partially denoised sequences expanded via masked diffusion and guided by multi-expert rollouts.
We propose Monte Carlo Tree Diffusion with Multi-Experts (MCTD-ME), which enhances protein generation by combining masked diffusion with an MCTS planner. An overview of the MCTD-ME workflow is shown in Fig 1. MCTD-ME structures diffusion as a tree, refining partially masked sequences to preserve scaffold integrity while systematically exploring uncertain regions. Search quality and diversity are further improved by incorporating multiple diffusion experts for diverse refinements and by using pLDDT-guided masking to target low-confidence residues, as illustrated in Fig 2. These design choices couple the planning strength of tree search with expert-aware diversification, yielding a principled framework for balancing exploration and exploitation in protein modeling. We present the algorithm in Algorithm 1 (deferred to Appendix A.1), with each rollout consisting of four key steps as following:
@@
To navigate the space of protein sequences under the reverse masked diffusion framework, we define a tree where each node corresponds to a complete sequence (with no masked tokens) at a particular reverse step, and edges represent one reverse transition obtained by applying masked discrete diffusion to a low-confidence subset of tokens. At each iteration, we employ a novel selection algorithm, $\text{P}\mathcal{H}\text{-UCT}\text{-ME}$—a Policy(P)-guided entropy($\mathcal{H}$)-augmented variant of UCT—designed to identify the most promising path, where ME denotes Multiple Experts. Specifically, from a parent node $s_{t}$ we choose the transition (i.e., reverse move to $s_{t-1}$ via a particular mask set) that maximizes the following objective:
$\text{P}\mathcal{H}\text{-UCT}\text{-ME}(s_{t})\;=\;\mathop{\rm arg\,max}_{a\in\mathcal{A}}\;\text{P}\mathcal{H}\text{-UCB}\text{-ME}\!\left((s_{t},a)\right),$
@@
$\text{P}\mathcal{H}\text{-UCB}\text{-ME}\!$ $\left((s_{t},a)\right)=Q(s_{t},a)+c_{p}\,\frac{\sqrt{\log N(s_{t})}}{1+N(s_{t},a)}$
$\;\pi_{\mathrm{cons},\tau}(a\mid s_{t})$ $\cdot\left({\underbrace{w_{\mathrm{ent}}\,\mathcal{U}_{\text{ent}}(s_{t},a)\;+\;w_{\mathrm{div}}\,\mathcal{U}_{\text{div}}(s_{t},a)}_{\text{$\text{P}\mathcal{H}$\text{-ME}~bonus (cached at expansion)}}}\right).$ (2)
Here, $Q(s_{t},a)$ is the current value estimate, $N(\cdot)$ are visit counts, $c_{p}>0$ controls exploration, and $w_{\mathrm{ent}},w_{\mathrm{div}}\!\geq\!0$ weight the predictive-uncertainty and diversity bonuses, respectively. In our diffusion setting, the action prior is given by the reverse diffusion model $p_{\phi}(s_{t-1}\!\mid\!s_{t})$ over masked edits, but—consistent with our implementation—we do not recompute its entropy during selection; instead we cache uncertainty statistics for each expanded child and reuse them during subsequent selections. We cache a temperature-controlled consensus prior $\pi_{\mathrm{cons},\tau}(a\mid s_{t})\!\in\![0,1]$ at expansion—large only when multiple experts assign high mass to the same action—so the UCB exploration is steered toward expert-agreed edits while preserving asymptotics. We define
$\pi_{\mathrm{cons},\tau}(a\mid s)=\frac{1}{Z_{\tau}}\exp\!\Big(\tfrac{1}{\tau}\sum_{e=1}^{E}\log p^{(e)}_{\phi}(a\mid s,\mathcal{M})\Big),$ (3)
$Z_{\tau}=\sum_{a^{\prime}\in\mathcal{A}(s)}\exp\!\Big(\tfrac{1}{\tau}\sum_{e=1}^{E}\log p^{(e)}_{\phi}(a^{\prime}\mid s,\mathcal{M})\Big),$ (4)
i.e., a temperatured geometric mean across experts.
[…]
$\mathcal{U}_{\text{ent}}(s_{t},a)=$ $\mathcal{H}\!\left(\frac{1}{E}\sum_{e=1}^{E}p_{\phi}^{(e)}(\,\cdot\,\mid s_{t-1},\mathcal{M})\right)$
$-\frac{1}{E}\sum_{e=1}^{E}\,\mathcal{H}\!\left(p_{\phi}^{(e)}(\,\cdot\,\mid s_{t-1},\mathcal{M})\right).$ (5)
Here $\mathcal{H}(p)=-\sum_{a}p(a)\log p(a)$ is Shannon entropy; the difference equals a BALD/Jensen–Shannon–style disagreement signal, isolating epistemic uncertainty.
@@
$\mathcal{U}_{\text{div}}(s_{t},a)$ $=\frac{1}{L}\sum_{i=1}^{L}\,\mathbf{1}\{y^{\text{child}}_{i}\neq y^{\text{parent}}_{i}\}.$ (6)
In our multi-model setting, sequences are batched and multi-channel (e.g., amino-acid tokens and structure tokens). We evaluate $\mathcal{U}_{\text{div}}$ over the task-relevant token subset: for folding, the difference is computed on structure tokens relative to the root structure; for inverse folding, on amino-acid tokens relative to the baseline sequence; and for motif scaffolding, on both channels (combined by concatenation or by averaging channel-wise normalized distances). When batching, we average the per-sequence $\mathcal{U}_{\text{div}}$ across the batch, thus the result remains in $[0,1]$. Both $\mathcal{U}_{\text{ent}}$ and $\mathcal{U}_{\text{div}}$ are computed once at expansion and cached for reuse during selection.
@@
Remark 1 (single-expert case). When E=1E{=}1, (5) vanishes ( $\mathcal{U}_{\text{ent}}{=}0$ ). In MCTD-1, we therefore use the predictive (Shannon) entropy of the single expert over the masked sites as the uncertainty bonus:
$\mathcal{U}_{\text{ent}}^{(1)}(s_{t},a)=\frac{1}{|\mathcal{M}|}\sum_{i\in\mathcal{M}}\mathcal{H}\!\big(p_{\phi}(x_{i}\mid s_{t-1},\mathcal{M})\big),$
which preserves the same units and interpretation (larger $\Rightarrow\,$ more uncertain edit). This value is computed at expansion and cached for reuse, akin to the entropy-guided UCB bonus used in ERP [23]. In the single-expert case, the multiplicative prior reduces to that expert’s $\pi_{\tau}(a\mid s_{t})$, ensuring a coherent reduction.